OpenStudy (anonymous):

which point lies in the line described by the equation below? y-3=2(x+5) a. 4,18 b.4,21 c.3,17 d.3,-3 e.-2,13 f.-1,9 when i solved it i got (0,13) but obviously this can't be correct as it isnt a choice can someone help me?

OpenStudy (campbell_st):

just substitute each pariir to check 1st pair, 4, 18 18 - 3 = 2(4 + 5) does 15 = 18...? same process for each pair... hopefully the crr correct answer isn't the last choice soes

OpenStudy (anonymous):

so would it be b (4,21) since 21-3=18 and 2(4+5)=18?

OpenStudy (campbell_st):

thats correct...